# Personalized Enhanced Federated Multi-View Clustering via Heat-Kernel Tensor Decomposition

### v2

@@ 6 Proposed Federated Tensorized Clustering
In this section, an introduction is provided to the proposed federated tensorized clustering framework, which integrates heat-kernel-enhanced multi-view clustering with tensor decomposition techniques. This integration enables the efficient management of high-dimensional multi-view data distributed across multiple clients. The central concept is to utilize tensor representations to capture the multi-modal structure of the data while leveraging heat-kernel based distance metrics to enhance clustering performance in a federated learning setting.
@@ 6.2 The Proposed FedHK-PARAFAC2 Local Objective Function
The FedHK-PARAFAC2 local objective function for client $\ell$ is defined as:
@@
The optimal updates for the tensorized federated heat-kernel enhanced multi-view clustering at client $\ell$ are given by:
Membership updates:
@@
Core tensor updates:
@@
View weight updates:
@@
Cluster center updates via PARAFAC2 decomposition:
@@
Factor matrix updates:
@@
By applying the method of Lagrange multipliers to the tensorized local objective in Eq. LABEL:eqn:tensor_local_obj and following similar derivation steps as in Theorem 1, we derive the update rules for membership, core tensor, factor matrices, and view weights. The key steps involve differentiating the objective with respect to each variable while enforcing the constraints, leading to the closed-form solutions presented in Eqs. 68 to 70. The alternating optimization approach ensures convergence to a local optimum under the same conditions outlined in Theorem 1.
First, we derive the membership update by differentiating the Lagrangian with respect to $\mu_{jk}^{(i)}$ and setting it to zero, leading to Eq. 68. Next, we derive the core tensor and factor matrix updates by fixing other variables and solving the resulting least-squares problems, yielding Eqs. 69 to 74. Finally, we derive the view weight update by differentiating with respect to $v_{h}^{(i)}$ and applying the normalization constraint, resulting in Eq. 70. Finally, we verify that these updates satisfy the KKT conditions and ensure convergence under the specified conditions.
@@ 6.3 The Proposed FedHK-Tucker Local Objective Function
By implementing Tucker decomposition in Eq. 63, the FedHK-Tucker local objective function for client $\ell$ is defined as:
@@
Theorem 3 (The FedHK-Tucker Update Rules).
The optimal updates for the tensorized federated heat-kernel enhanced multi-view clustering with Tucker decomposition at client $\ell$ are given by: Membership updates:
@@
Core tensor updates:
@@
View weight updates:
@@
Cluster center updates via Tucker decomposition:
@@
Factor matrix updates:
@@
The proof follows similar steps as in Theorem 2, with adjustments for the Tucker decomposition structure. By applying the method of Lagrange multipliers to the Tucker-based local objective in Eq. 75 and differentiating with respect to each variable while enforcing the constraints, we derive the closed-form solutions presented in Eqs. 79 to 81. The alternating optimization approach ensures convergence to a local optimum under the same conditions outlined in Theorem 1. ∎
@@ 6.4 Local Clustering Algorithms
Provided below the differences between the local clustering algorithms for E-FKMVC and the tensorized FedHK-PARAFAC2/FedHK-Tucker methods. Table 1 summarizes the key distinctions.
Table 1: Comparison of Local Clustering Algorithms
Aspect E-FKMVC FedHK-PARAFAC2/FedHK-Tucker
Data Structure Multi-view matrices Multi-view tensors
Distance Metric FKED (Eq. 4) TKED (Eq. 60)
Cluster Center Representation Direct centers $A^{h}$ Decomposed tensors $\mathcal{G}_{[\ell]},\mathbf{P}_{[\ell]},\mathbf{Q}_{[\ell]},\mathbf{R}_{[\ell]}$ or $\mathcal{G}_{[\ell]},\mathbf{U}_{[\ell]},\mathbf{V}_{[\ell]},\mathbf{W}_{[\ell]}$
Update Rules Eqs. 26 - 28 Eqs. 68 - 70 or 79 - 81
Optimization Method Direct updates Alternating Least Squares (ALS) for tensor factors
@@
The PARAFAC2 and Tucker decompositions in E-FKMVC enable efficient representation and updating of cluster centers in high-dimensional multi-view spaces, facilitating effective clustering in federated settings. For both tensorized methods, the local clustering algorithms involve iterative updates of memberships, tensor factors, and view weights until convergence. Table 2 summarizes the key differences between the FedHK-PARAFAC2 and FedHK-Tucker methods.
Table 2: Comparison of FedHK-PARAFAC2 and FedHK-Tucker Methods
@@
Algorithm 2 Tensorized Federated Heat-Kernel Enhanced Multi-View Clustering (Local FedHK-PARAFAC2)
@@
Input: Local tensor data $\mathcal{X}_{[\ell]}$, global tensor factors $\{\mathcal{G}_{g},\mathbf{P}_{g},\mathbf{Q}_{g},\mathbf{R}_{g}\}$, personalization parameters $\{\lambda_{[\ell]},\rho_{[\ell]}\}$, $g$ is the global model index
Initialize:
$\mathcal{G}_{[\ell]}^{(0)}\leftarrow\lambda_{[\ell]}\mathcal{G}_{g}+(1-\lambda_{[\ell]})\mathcal{G}_{[\ell]}$
$\mathbf{P}_{[\ell]}^{(0)},\mathbf{Q}_{[\ell]}^{(0)},\mathbf{R}_{[\ell]}^{(0)}\leftarrow$ similarly personalized initialization
$v_{[\ell]}^{(0)}\leftarrow\frac{1}{s(\ell)}\mathbf{1}$
$t\leftarrow 0$
repeat
  $t\leftarrow t+1$
  // Compute tensorized heat-kernel coefficients
  Calculate $\delta_{[\ell]ij}^{h}$ using Eq. 61
  // Reconstruct cluster centers from decomposition
  $\mathcal{A}_{[\ell]}\leftarrow\mathcal{G}_{[\ell]}\times_{1}\mathbf{P}_{[\ell]}\times_{2}\mathbf{Q}_{[\ell]}\times_{3}\mathbf{R}_{[\ell]}$
  // Update memberships
  Update $U_{[\ell]}^{*^{(t)}}$ using Eq. 68
  // Update tensor factors using ALS
  Update $\mathcal{G}_{[\ell]}^{(t)}$ using Eq. 69
  Update $\mathbf{P}_{[\ell]}^{(t)}$ using Eq. 72
  Update $\mathbf{Q}_{[\ell]}^{(t)}$ using Eq. 73
  Update $\mathbf{R}_{[\ell]}^{(t)}$ using Eq. 74
  // Update view weights
  Update $v_{[\ell]}^{(t)}$ using Eq. 70
  // Check convergence
  Compute $J_{[\ell]}^{tensor}$ using Eq. 64
until convergence or max iterations
return $U_{[\ell]}^{*^{(t)}},\mathcal{G}_{[\ell]}^{(t)},\mathbf{P}_{[\ell]}^{(t)},\mathbf{Q}_{[\ell]}^{(t)},\mathbf{R}_{[\ell]}^{(t)},v_{[\ell]}^{(t)}$
Algorithm 3 Tensorized Federated Heat-Kernel Enhanced Multi-View Clustering (Local FedHK-Tucker)
@@
Input: Local tensor data $\mathcal{X}_{[\ell]}$, global tensor factors $\{\mathcal{G}_{g},\mathbf{U}_{g},\mathbf{V}_{g},\mathbf{W}_{g}\}$, personalization parameters $\{\lambda_{[\ell]},\rho_{[\ell]}\}$, $g$ is the global model index
Initialize:
$\mathcal{G}_{[\ell]}^{(0)}\leftarrow\lambda_{[\ell]}\mathcal{G}_{g}+(1-\lambda_{[\ell]})\mathcal{G}_{[\ell]}$
$\mathbf{U}_{[\ell]}^{(0)},\mathbf{V}_{[\ell]}^{(0)},\mathbf{W}_{[\ell]}^{(0)}\leftarrow$ similarly personalized initialization
$v_{[\ell]}^{(0)}\leftarrow\frac{1}{s(\ell)}\mathbf{1}$
$t\leftarrow 0$
repeat
  $t\leftarrow t+1$
  // Compute tensorized heat-kernel coefficients
  Calculate $\delta_{[\ell]ij}^{h}$ using Eq. 61
  // Reconstruct cluster centers from decomposition
  $\mathcal{A}_{[\ell]}\leftarrow\mathcal{G}_{[\ell]}\times_{1}\mathbf{U}_{[\ell]}\times_{2}\mathbf{V}_{[\ell]}\times_{3}\mathbf{W}_{[\ell]}$
  // Update memberships
  Update $U_{[\ell]}^{*^{(t)}}$ using Eq. 79
  // Update tensor factors using ALS
  Update $\mathcal{G}_{[\ell]}^{(t)}$ using Eq. 80
  Update $\mathbf{U}_{[\ell]}^{(t)}$ using Eq. 83
  Update $\mathbf{V}_{[\ell]}^{(t)}$ using Eq. 84
  Update $\mathbf{W}_{[\ell]}^{(t)}$ using Eq. 85
  // Update view weights
  Update $v_{[\ell]}^{(t)}$ using Eq. 81
  // Check convergence
  Compute $J_{[\ell]}^{tensor}$ using Eq. 75
until convergence or max iterations
return $U_{[\ell]}^{*^{(t)}},\mathcal{G}_{[\ell]}^{(t)},\mathbf{U}_{[\ell]}^{(t)},\mathbf{V}_{[\ell]}^{(t)},\mathbf{W}_{[\ell]}^{(t)},v_{[\ell]}^{(t)}$

### v3

@@ 4 Proposed Federated Tensorized Clustering
In this section, an introduction is provided to the proposed federated tensorized clustering framework, which integrates efficient kernel multi-view clustering with tensor decomposition techniques. This integration enables the efficient management of high-dimensional multi-view data distributed across multiple clients. The central concept is to utilize tensor representations to capture the multi-modal structure of the data while leveraging heat-kernel based distance metrics to enhance clustering performance in a federated learning setting.
@@ 4.2 The Proposed FedHK-PARAFAC2 Local Objective Function
Based on E-FKMVC and the tensorized representations, we formulate the local objective function for the proposed Federated Heat-Kernel Enhanced Clustering with PARAFAC2 decomposition (FedHK-PARAFAC2) at client $\ell$. The objective function aims to minimize the weighted sum of tensorized kernel Euclidean distances between data points and their corresponding cluster centers, while incorporating membership degrees and view weights. The local objective function is defined as follows:
@@
The optimal updates for the tensorized efficient federated kernel multi-view clustering at client $\ell$ are given by the following iterative formulas. These update rules ensure that each component of the model—memberships, tensor factors, and view weights—is optimized while holding the other components fixed, following the alternating optimization strategy.
Membership updates: The membership matrix is a crucial component in determining the extent to which each data point aligns with a specific cluster. The update rule ascribes elevated membership values to clusters whose centers are proximate to the data point (in the heat-kernel weighted sense).
@@
Here, $\mu_{[\ell]ik}^{*}$ represents the degree to which data point $i$ belongs to cluster $k$. The negative exponent 1/(m1)-1/(m-1) ensures that smaller distances (indicating closer proximity to a cluster center) result in larger membership values. The denominator normalizes these values so that the memberships for each data point sum to one across all clusters.
Core tensor updates: The core tensor $\mathcal{G}_{[\ell]}$ captures the essential interactions between the different modes (clusters, features, and views) in a compressed form. It is updated by solving a weighted least-squares problem that minimizes the reconstruction error between the original data and the decomposed cluster centers:
@@
This optimization finds the core tensor that best represents the cluster structure when combined with the factor matrices. The weights $(\mu_{[\ell]ik}^{*})^{m}$ and $(v_{[\ell]h})^{\alpha}$ ensure that more confident cluster assignments and more important views contribute more strongly to the update.
View weight updates: The view weights $v_{[\ell]h}$ quantify the relative importance of each data view in the clustering process. Views that produce more coherent clustering results (i.e., tighter clusters with smaller within-cluster distances) receive higher weights:
@@
The structure mirrors that of the membership update: views with smaller total weighted distances receive larger weights, and the denominator ensures that all view weights sum to one.
Cluster center updates via PARAFAC2 decomposition: The cluster center tensor is reconstructed from its PARAFAC2 decomposition components using the $n$-mode product operations:
@@
This compact representation expresses the full cluster center tensor through the smaller core tensor $\mathcal{G}_{[\ell]}$ and three factor matrices $\mathbf{P}_{[\ell]}$, $\mathbf{Q}_{[\ell]}$, and $\mathbf{R}_{[\ell]}$, corresponding to the cluster, feature, and view modes held by client $\ell$, respectively. This decomposition significantly reduces storage and communication requirements in the federated setting.
Factor matrix updates: Each factor matrix is updated by solving a least-squares problem while holding all other components fixed. These updates are performed using the Alternating Least Squares (ALS) procedure:
@@
The factor matrix $\mathbf{P}_{[\ell]}$ captures the cluster mode structure, $\mathbf{Q}_{[\ell]}$ captures the feature mode structure, and $\mathbf{R}_{[\ell]}$ captures the view mode structure. By iteratively updating each factor while fixing the others, the ALS procedure converges to a locally optimal decomposition that accurately represents the cluster centers.
@@
By applying the method of Lagrange multipliers to the tensorized local objective in Eq. 71 and following similar derivation steps as in Theorem 1, we derive the update rules for membership, core tensor, factor matrices, and view weights. The key steps involve differentiating the objective with respect to each variable while enforcing the constraints, leading to the closed-form solutions presented in Eqs. 75 to 77. The alternating optimization approach ensures convergence to a local optimum under the same conditions outlined in Theorem 1.
First, we derive the membership update by differentiating the Lagrangian with respect to $\mu_{[\ell]ik}^{*}$, leading to Eq. 75. Next, we update the core tensor $\mathcal{G}_{[\ell]}$ by solving the weighted least-squares problem in Eq. 76. The view weights are updated by differentiating with respect to $v_{[\ell]h}$, resulting in Eq. 77. Finally, the cluster center tensor is reconstructed using the PARAFAC2 decomposition in Eq. 78, and the factor matrices are updated via ALS as shown in Eqs. 79 to 81. ∎
@@ 4.3 The Proposed FedHK-Tucker Local Objective Function
Similar to the FedHK-PARAFAC2 framework, we formulate the local objective function for the proposed Federated Heat-Kernel Enhanced Clustering with Tucker decomposition (FedHK-Tucker) at client $\ell$. The objective function aims to minimize the weighted sum of tensorized kernel Euclidean distances between data points and their corresponding cluster centers, while incorporating membership degrees and view weights. The local objective function is defined as follows:
@@
Theorem 3 (Update Rules for FedHK-Tucker).
For the tensorized efficient federated kernel multi-view clustering framework employing Tucker decomposition at client $\ell$, the following iterative update formulas yield optimal solutions. The alternating optimization strategy ensures that each model component—memberships, tensor factors, and view weights—is optimized independently while other components remain fixed.
Membership Matrix Updates: The membership coefficients determine the degree of association between data points and clusters. Higher membership values are assigned to clusters with centers in closer proximity to the data point, as measured by the heat-kernel weighted distance:
@@
The coefficient $\mu_{[\ell]ik}^{*}$ quantifies the membership degree of data point $i$ in cluster $k$ held by client $\ell$. The exponent 1/(m1)-1/(m-1) establishes an inverse relationship between distance and membership, whereby smaller distances yield larger membership values. Normalization through the denominator guarantees that memberships across all clusters sum to unity for each data point.
Core Tensor Updates: The core tensor $\mathcal{G}_{[\ell]}$ encodes the fundamental interactions among clusters, features, and views in a compressed representation. Its update involves minimizing a weighted least-squares objective that quantifies the reconstruction error:
@@
This formulation identifies the core tensor that optimally captures the cluster structure in conjunction with the factor matrices. The weighting terms $(\mu_{[\ell]ik}^{*})^{m}$ and $(v_{[\ell]h})^{\alpha}$ amplify contributions from high-confidence cluster assignments and significant views.
View Weight Updates: The view weights $v_{[\ell]h}$ encode the relative significance of each data view within the clustering procedure. Views producing more coherent clusters, characterized by smaller within-cluster distances, receive elevated weights:
@@
This formulation parallels the membership update structure: views exhibiting smaller aggregate weighted distances obtain larger weights, with the denominator ensuring unit summation across all view weights.
Cluster Center Reconstruction via Tucker Decomposition: The cluster center tensor is assembled from its Tucker decomposition constituents through successive $n$-mode products:
@@
This factorized representation expresses the complete cluster center tensor via the compact core tensor $\mathcal{G}_{[\ell]}$ combined with three factor matrices: $\mathbf{U}_{[\ell]}$ for clusters, $\mathbf{V}_{[\ell]}$ for features, and $\mathbf{W}_{[\ell]}$ for views at client $\ell$. This decomposition substantially reduces storage and communication overhead in federated environments.
Factor Matrix Updates: Individual factor matrices are refined through least-squares optimization with all remaining components held constant. The Alternating Least Squares (ALS) procedure governs these updates:
@@
The matrix $\mathbf{U}_{[\ell]}$ models cluster mode structure, $\mathbf{V}_{[\ell]}$ represents feature mode structure, and $\mathbf{W}_{[\ell]}$ encodes view mode structure. Through cyclic updates of each factor with others fixed, the ALS procedure attains a locally optimal decomposition that faithfully represents the cluster centers.
@@
The proof follows similar steps as in Theorem 2, with adjustments for the Tucker decomposition structure. By applying the method of Lagrange multipliers to the Tucker-based local objective in Eq. 82 and differentiating with respect to each variable while enforcing the constraints, we derive the closed-form solutions presented in Eqs. 86 to 88. The alternating optimization approach ensures convergence to a local optimum under the same conditions outlined in Theorem 1. ∎
@@ 4.4 Local Clustering Algorithms
Provided below the differences between the local clustering algorithms for E-FKMVC and the tensorized FedHK-PARAFAC2/FedHK-Tucker methods. Table 2 summarizes the key distinctions. While both approaches aim to achieve effective multi-view clustering in a federated setting, they differ significantly in their data representation, distance metrics, cluster center representations, and update mechanisms. Table 3 further contrasts the FedHK-PARAFAC2 and FedHK-Tucker methods, highlighting their unique aspects related to tensor decomposition and update rules. For concise summary of the tensor decomposition employed in both methods, please refer to Table 4.
Table 2: Comparison of Local Clustering Algorithms for E-FKMVC and Tensorized FedHK Methods. The tensorized methods include both FedHK-PARAFAC2 and FedHK-Tucker.
Aspect E-FKMVC FedHK-PARAFAC2/FedHK-Tucker Data Structure Multi-view matrices Multi-view tensors Distance Metric FKED (Eq. 4) TKED (Eq. 67) Cluster Center Representation Direct centers $A^{h}$ Decomposed tensors $\mathcal{G}_{[\ell]},\mathbf{P}_{[\ell]},\mathbf{Q}_{[\ell]},\mathbf{R}_{[\ell]}$ or $\mathcal{G}_{[\ell]},\mathbf{U}_{[\ell]},\mathbf{V}_{[\ell]},\mathbf{W}_{[\ell]}$ Update Rules Eqs. 29 - 31 Eqs. 75 - 77 or 86 - 88 Optimization Method Direct updates Alternating Least Squares (ALS) for tensor factors
@@
Table 3: Comparison of FedHK-PARAFAC2 and FedHK-Tucker Methods in terms of Key Aspects to Tensor Decomposition and Update Rules.
@@
Table 4: Comparison of PARAFAC2 and Tucker Decomposition Methods for Cluster Center Representation in FedHK-PARAFAC2 and FedHK-Tucker.
Aspect PARAFAC2 Tucker Decomposition Type PARAFAC2 (Parallel Factor Analysis 2) Tucker Decomposition Cluster Center Representation $\mathcal{A}_{[\ell]}=\mathcal{G}_{[\ell]}\times_{1}\mathbf{P}_{[\ell]}\times_{2}\mathbf{Q}_{[\ell]}\times_{3}\mathbf{R}_{[\ell]}$ $\mathcal{A}_{[\ell]}=\mathcal{G}_{[\ell]}\times_{1}\mathbf{U}_{[\ell]}\times_{2}\mathbf{V}_{[\ell]}\times_{3}\mathbf{W}_{[\ell]}$ Key Components Core tensor $\mathcal{G}_{[\ell]}$ Factor matrices $\mathbf{P}_{[\ell]},\mathbf{Q}_{[\ell]},\mathbf{R}_{[\ell]}$ Core tensor $\mathcal{G}_{[\ell]}$ Factor matrices $\mathbf{U}_{[\ell]},\mathbf{V}_{[\ell]},\mathbf{W}_{[\ell]}$ Salient Features Handles varying dimensions across one mode Maintains consistent factors in other modes Suitable for heterogeneous client data Unique decomposition under mild conditions Flexible multilinear ranks $(R_{1},R_{2},R_{3})$ Higher compression ratio Captures complex multilinear interactions More flexible but non-unique decomposition Optimization Method ALS for PARAFAC2 factors (Eqs. 7681) ALS for Tucker factors (Eqs. 8792) Storage Complexity $\mathcal{O}(r_{1}r_{2}r_{3}+cr_{1}+Dr_{2}+sr_{3})$ $\mathcal{O}(R_{1}R_{2}R_{3}+cR_{1}+DR_{2}+sR_{3})$ Best Suited For Data with consistent structure across views but varying sample dimensions Data with complex interactions requiring flexible rank selection per mode
As can be seen in Table 4, PARAFAC2 is advantageous for handling heterogeneous client data with varying dimensions across one mode while maintaining consistent factors in other modes. In contrast, Tucker decomposition offers greater flexibility through multilinear ranks, allowing for higher compression ratios and the ability to capture complex interactions among modes. The choice between the two methods depends on the specific characteristics of the multi-view data and the clustering requirements in the federated setting.
@@
Algorithm 2 FedHK-PARAFAC2 Local Clustering Algorithm
@@
0: Local tensor data $\mathcal{X}_{[\ell]}\in\mathbb{R}^{n(\ell)\times D_{[\ell]}\times s(\ell)}$, global tensor factors $\{\mathcal{G}_{g},\mathbf{P}_{g},\mathbf{Q}_{g},\mathbf{R}_{g}\}$, personalization parameters $\{\lambda_{[\ell]},\rho_{[\ell]}\}$, fuzzifier $m>1$, view exponent $\alpha>1$, convergence threshold $\epsilon>0$, max iterations $T_{max}$
0: Local membership matrix $U_{[\ell]}^{*}$, tensor factors $\{\mathcal{G}_{[\ell]},\mathbf{P}_{[\ell]},\mathbf{Q}_{[\ell]},\mathbf{R}_{[\ell]}\}$, view weights $v_{[\ell]}$
1:/* Phase 1: Personalized Initialization */
[…]
9:/* Phase 2: Iterative Optimization */
10:repeat
11:  $t\leftarrow t+1$
12:  /* Step 2.1: Heat-Kernel Coefficient Computation */
13:  for $h=1$ to $s(\ell)$ do
14:   for $j=1$ to $d_{h}^{(\ell)}$ do
15:    $\delta_{[\ell]ij}^{h}\leftarrow\frac{\mathcal{X}_{[\ell],i,j}^{h}-\min_{i^{\prime}}\mathcal{X}_{[\ell],i^{\prime},j}^{h}}{\max_{i^{\prime}}\mathcal{X}_{[\ell],i^{\prime},j}^{h}-\min_{i^{\prime}}\mathcal{X}_{[\ell],i^{\prime},j}^{h}+\epsilon}$ {Eq. 68}
16:   end for
17:  end for
18:  /* Step 2.2: Cluster Center Reconstruction */
19:  $\mathcal{A}_{[\ell]}^{(t)}\leftarrow\mathcal{G}_{[\ell]}^{(t-1)}\times_{1}\mathbf{P}_{[\ell]}^{(t-1)}\times_{2}\mathbf{Q}_{[\ell]}^{(t-1)}\times_{3}\mathbf{R}_{[\ell]}^{(t-1)}$ {Eq. 78}
20:  /* Step 2.3: Membership Matrix Update */
21:  for $i=1$ to $n(\ell)$ do
22:   for $k=1$ to $c(\ell)$ do
23:    Compute $\mu_{[\ell]ik}^{*^{(t)}}$ using Eq. 75
24:   end for
25:  end for
26:  /* Step 2.4: Tensor Factor Updates via ALS */
27:  Update $\mathcal{G}_{[\ell]}^{(t)}$ by solving Eq. 76
28:  Update $\mathbf{P}_{[\ell]}^{(t)}$ by solving Eq. 79
29:  Update $\mathbf{Q}_{[\ell]}^{(t)}$ by solving Eq. 80
30:  Update $\mathbf{R}_{[\ell]}^{(t)}$ by solving Eq. 81
31:  /* Step 2.5: View Weight Update */
32:  for $h=1$ to $s(\ell)$ do
33:   Compute $v_{[\ell]h}^{(t)}$ using Eq. 77
34:  end for
35:  /* Step 2.6: Convergence Check */
36:  Compute $J_{[\ell]}^{(t)}$ using Eq. 71
37:  $\Delta J\leftarrow|J_{[\ell]}^{(t)}-J_{[\ell]}^{(t-1)}|/|J_{[\ell]}^{(t-1)}|$
38:until $\Delta J<\epsilon$ or $t\geq T_{max}$
39:/* Phase 3: Return Results */
40:return $U_{[\ell]}^{*^{(t)}},\mathcal{G}_{[\ell]}^{(t)},\mathbf{P}_{[\ell]}^{(t)},\mathbf{Q}_{[\ell]}^{(t)},\mathbf{R}_{[\ell]}^{(t)},v_{[\ell]}^{(t)}$
Algorithm 3 FedHK-Tucker Local Clustering Algorithm
@@
0: Local tensor data $\mathcal{X}_{[\ell]}\in\mathbb{R}^{n(\ell)\times D_{[\ell]}\times s(\ell)}$, global tensor factors $\{\mathcal{G}_{g},\mathbf{U}_{g},\mathbf{V}_{g},\mathbf{W}_{g}\}$, personalization parameters $\{\lambda_{[\ell]},\rho_{[\ell]}\}$, fuzzifier $m>1$, view exponent $\alpha>1$, convergence threshold $\epsilon>0$, max iterations $T_{max}$
0: Local membership matrix $U_{[\ell]}^{*}$, tensor factors $\{\mathcal{G}_{[\ell]},\mathbf{U}_{[\ell]},\mathbf{V}_{[\ell]},\mathbf{W}_{[\ell]}\}$, view weights $v_{[\ell]}$
1:/* Phase 1: Personalized Initialization */
[…]
9:/* Phase 2: Iterative Optimization */
10:repeat
11:  $t\leftarrow t+1$
12:  /* Step 2.1: Heat-Kernel Coefficient Computation */
13:  for $h=1$ to $s(\ell)$ do
14:   for $j=1$ to $d_{h}^{(\ell)}$ do
15:    $\delta_{[\ell]ij}^{h}\leftarrow\frac{\mathcal{X}_{[\ell],i,j}^{h}-\min_{i^{\prime}}\mathcal{X}_{[\ell],i^{\prime},j}^{h}}{\max_{i^{\prime}}\mathcal{X}_{[\ell],i^{\prime},j}^{h}-\min_{i^{\prime}}\mathcal{X}_{[\ell],i^{\prime},j}^{h}+\epsilon}$ {Eq. 68}
16:   end for
17:  end for
18:  /* Step 2.2: Cluster Center Reconstruction */
19:  $\mathcal{A}_{[\ell]}^{(t)}\leftarrow\mathcal{G}_{[\ell]}^{(t-1)}\times_{1}\mathbf{U}_{[\ell]}^{(t-1)}\times_{2}\mathbf{V}_{[\ell]}^{(t-1)}\times_{3}\mathbf{W}_{[\ell]}^{(t-1)}$ {Eq. 89}
20:  /* Step 2.3: Membership Matrix Update */
21:  for $i=1$ to $n(\ell)$ do
22:   for $k=1$ to $c(\ell)$ do
23:    Compute $\mu_{[\ell]ik}^{*^{(t)}}$ using Eq. 86
24:   end for
25:  end for
26:  /* Step 2.4: Tensor Factor Updates via ALS */
27:  Update $\mathcal{G}_{[\ell]}^{(t)}$ by solving Eq. 87
28:  Update $\mathbf{U}_{[\ell]}^{(t)}$ by solving Eq. 90
29:  Update $\mathbf{V}_{[\ell]}^{(t)}$ by solving Eq. 91
30:  Update $\mathbf{W}_{[\ell]}^{(t)}$ by solving Eq. 92
31:  /* Step 2.5: View Weight Update */
32:  for $h=1$ to $s(\ell)$ do
33:   Compute $v_{[\ell]h}^{(t)}$ using Eq. 88
34:  end for
35:  /* Step 2.6: Convergence Check */
36:  Compute $J_{[\ell]}^{(t)}$ using Eq. 82
37:  $\Delta J\leftarrow|J_{[\ell]}^{(t)}-J_{[\ell]}^{(t-1)}|/|J_{[\ell]}^{(t-1)}|$
38:until $\Delta J<\epsilon$ or $t\geq T_{max}$
39:/* Phase 3: Return Results */
40:return $U_{[\ell]}^{*^{(t)}},\mathcal{G}_{[\ell]}^{(t)},\mathbf{U}_{[\ell]}^{(t)},\mathbf{V}_{[\ell]}^{(t)},\mathbf{W}_{[\ell]}^{(t)},v_{[\ell]}^{(t)}$